# DianJin-OCR-R1: Enhancing OCR Capabilities via a Reasoning-and-Tool Interleaved Vision-Language Model

### v1

@@ Abstract
Recent advances in large vision-language models (LVLMs) have enabled a new paradigm of end-to-end document image parsing, excelling in Optical Character Recognition (OCR) tasks such as text, table, and formula recognition. However, generative LVLMs, similarly to large language models (LLMs), are prone to hallucinations——generating words that do not exist in input images. Furthermore, LVLMs are designed for general purposes and tend to be less effective on OCR tasks compared to expert models that are trained on domain-specific datasets. In this paper, we propose DianJin-OCR-R1, a reasoning-enhanced framework designed to address these limitations through training reasoning-and-tool interleaved VLMs. Given a recognition instruction, our DianJin-OCR-R1 model first recognizes the content in the input image by its own OCR capabilities, and then calls other tools (i.e., other expert models) to obtain their results as references, finally ”looks again” the image and rethinks about the reasoning process to provide the final recognized content. Since architectures of expert models are tailored for specific OCR tasks, which makes them less prone to hallucinations, their results can help VLMs mitigate hallucinations. Additionally, expert models are typically smaller in scale and easy to iterate, enabling performance improvements for VLMs at a lower cost. We evaluate our model on ReST (Yu et al., 2023) and OmniDocBench (Ouyang et al., 2025), and experimental results show that our DianJin-OCR-R1 models consistently outperform their non-reasoning counterparts and expert OCR models, which proves the effectiveness of our method.
@@ 4.1 Experimental Setups
Datasets size $i_{token}$ $q_{token}$ $r_{token}$ $a_{token}$
@@ 4.2 Experimental Results
  • On the three test sets, our DianJin-OCR-R1 models significantly outperform the base model (Qwen2.5-VL-7B-Instruct). For instance, our DianJin-OCR-R1 model trained by SFT improves accuracy from 0.527 to 0.722 on seal recognition task, TEDS from 0.858 to 0.895 on table recognition task, and CDM from 0.915 to 0.977 on formula recognition task.
@@
  • We achieve best results with the model using only thousands of training samples. Under the same data conditions, the models trained by RFT are generally superior to those trained by SFT, indicating that RFT is more advantageous.
@@ 4.3 Ablation Study
  • Baseline vs Baseline with Tools. Directly supplying the outputs of two tools as references to the baseline Qwen2.5-VL-7B-Instruct significantly boosts its performance (e.g., increasing accuracy from 0.551 to 0.708 on seal recognition tasks), which demonstrates that leveraging expert models can effectively enhance the capabilities of VLMs on OCR tasks.
  • Fine-tuning directly vs Fine-tuning with Reasoning and Tools. We conducted direct SFT and RFT on the base model, using only the image $i$ and question $a$ as inputs and the answer $a$ as output, omitting the reasoning chain $r$. The results indicate that models incorporating reasoning chains and tools consistently achieve superior performance, highlighting the successful integration of reasoning and tools.
  • Baseline with tools vs Fine-tuning with Reasoning and Tools. Comparing to the baseline with two tools, models trained by SFT with reasoning and tools perform better, which demonstrates the effectiveness of distillation method.
@@
Table 5: Ablation study
Furthermore, we also replace the tools to conduct more in-depth exploration. For formula recognition task, we replace MonkeyOCR-3B and PP-FormulaNet_plus-M with GOT and PP-FormulaNet_plus-S. As shown in Table 6, the performance using MonkeyOCR-3B and PP-FormulaNet_plus-M are better than using GOT and PP-FormulaNet_plus-S, which demonstrates that optimizing a expert OCR model alone whose costs are lower can improve the OCR capabilities of our model.
@@
Tool2: PP-FormulaNet_plus-M 0.966 0.730 0.197
DianJin-OCR-R1 (w tools) 0.979 0.770 0.189
@@ 5 Conclusion
We have presented DianJin-OCR-R1, a reasoning-augmented framework for large vision-language models on OCR tasks. This framework fosters a straightforward reasoning paradigm in VLMs, enabling a more transparent and verifiable reasoning process on complex document OCR tasks. Through extensive experiments on diverse benchmarks, we demonstrate that reasoning-aware training significantly improves both performance.

### v2

@@ Abstract
Recent advances in large vision-language models (LVLMs) have enabled a new paradigm of end-to-end document image parsing, excelling in Optical Character Recognition (OCR) tasks such as text, table, and formula recognition. However, generative LVLMs, similarly to large language models (LLMs), are prone to hallucinations——generating words that do not exist in input images. Furthermore, LVLMs are designed for general purposes and tend to be less effective on OCR tasks compared to expert models that are trained on domain-specific datasets. In this paper, we propose DianJin-OCR-R1, a reasoning-enhanced framework designed to address these limitations through training reasoning-and-tool interleaved VLMs. Given a recognition instruction, our DianJin-OCR-R1 model first recognizes the content in the input image by its own OCR capabilities, and then calls other tools (i.e., other expert models) to obtain their results as references, finally ”looks again” the image and rethinks about the reasoning process to provide the final recognized content. Since architectures of expert models are tailored for specific OCR tasks, which makes them less prone to hallucinations, their results can help VLMs mitigate hallucinations. We evaluate our model on ReST and OmniDocBench, and experimental results show that our DianJin-OCR-R1 models consistently outperform their non-reasoning counterparts and expert OCR models, which proves the effectiveness of our method. Additionally, the results indicate that enhancing expert models, which are typically small and easy to iterate, enable performance improvements for VLMs.
@@ 4.1 Experimental Setups
Datasets size $q_{token}$ $i_{token}$ $r_{token}$ $a_{token}$
@@ 4.2 Experimental Results
  • Our DianJin-OCR-R1 models significantly outperform the base model (Qwen2.5-VL-7B-Instruct) on the three test sets. For instance, our DianJin-OCR-R1 model trained by SFT improves accuracy from 0.527 to 0.722 on seal recognition task, TEDS from 0.858 to 0.895 on table recognition task, and CDM from 0.915 to 0.977 on formula recognition task.
@@
  • Our DianJin-OCR-R1 models achieve best results with the model using only thousands of training samples. Under the same data conditions, the models trained by RFT are generally superior to those trained by SFT, indicating that RFT is more advantageous.
@@ 4.3 Ablation Study
  • Baseline vs Baseline with Tools. Directly supplying the outputs of two tools as references to the baseline Qwen2.5-VL-7B-Instruct significantly boosts its performance (e.g., increasing accuracy from 0.527 to 0.560 on the seal recognition task), which demonstrates that leveraging expert models can effectively enhance the capabilities of VLMs on OCR tasks.
  • Fine-tuning directly vs Fine-tuning with Reasoning and Tools. Comparing to models trained by SFT and RFT directly (only using the answer $a$ as output and omitting the reasoning chain $r$), our models incorporating reasoning chains and tools consistently achieve superior performance, which highlights the successful integration of reasoning and tools.
  • Baseline with tools vs Fine-tuning with Reasoning and Tools. Comparing to the baseline with two tools, our models trained by SFT with reasoning and tools perform better, which demonstrates the effectiveness of distillation method.
@@
Table 5: Performance comparison of models using different training strategies.
Furthermore, we also replace the tools to conduct more in-depth exploration. For formula recognition task, we replace MonkeyOCR-3B and PP-FormulaNet_plus-M with GOT and PP-FormulaNet_plus-S. As shown in Table 6, the performance using MonkeyOCR-3B and PP-FormulaNet_plus-M are better than using GOT and PP-FormulaNet_plus-S, which demonstrates that optimizing a expert OCR model alone whose costs are lower can improve the OCR capabilities of our models.
@@
Tool2: PP-FormulaNet_plus-M 0.966 0.730 0.198
DianJin-OCR-R1 (w tools) 0.977 0.770 0.188
@@ 5 Conclusion
We have presented DianJin-OCR-R1, a reasoning-augmented framework for large vision-language models on OCR tasks. This framework fosters a straightforward reasoning paradigm in VLMs, enabling a more transparent and verifiable reasoning process on complex document OCR tasks. Through extensive experiments on diverse benchmarks, we demonstrate that reasoning-aware training significantly improves performance.